OpenStudy (anonymous):

If α and B are the measures of two first quadrants angles and sinα=4/5 and sin B= 5/13 find sin (α+B)

sam (.sam.):

Find sinα=4/5 and sin B= 5/13

sam (.sam.):

sinα=4/5 α=53.13, 126.87 ---------------------- sin B= 5/13 B=22.62, 157.38 then sin (α+B) sin (53.13+22.62)................(1) sin(126.87+157.38)..............(2)

sam (.sam.):

just ask if u didn't get it :)